OpenStudy (anonymous):

I decided to learn the mathematica programming langauge recently, and have been wondering: What is the difference between "=" and"=="

OpenStudy (anonymous):

when you a=5, it means you assign 5 to a. When you say a^2 - 5 a + 3 ==0, it means what are the values of a that satisfies the equation a^2 - 5 a + 3 =0

OpenStudy (anonymous):

If you issue the command Solve[a^2 - 5 a + 3 ==0,a] \[ \left\{\left\{a\to \frac{1}{2} \left(5-\sqrt{13}\right)\right\},\left\{ a\to \frac{1}{2} \left(5+\sqrt{13}\right)\right\}\right\} \]

OpenStudy (anonymous):

Ok, this makes since, thanks.